# Unlearning Inversion Attacks for Graph Neural Networks

### v1

@@ 5.2 The Proposed TrendAttack
This formulation is flexible and covers several existing MIA methods. For example, when $\phi([\mathbf{x}_{i}~{}||~{}\mathbf{p}_{i}],[\mathbf{x}_{j}~{}||~{}\mathbf{p}_{j}])=\mathbf{h}^{\top}\cdot\mathrm{MLP}(\mathbf{p}_{i},\mathbf{p}_{j})$, the model recovers MIA-GNN olatunji2021membership . It can also recover the StealLink attack he2021stealing by incorporating manually defined similarity features into $\phi$. This flexible structure allows our attack model to incorporate any existing MIA method as the backbone, ensuring it performs at least as well as prior approaches.
@@
In addition, as indicated by Claim 5.1 and Claim 5.2, a key challenge in graph unlearning inversion is to distinguish nodes associated with unlearned edges from others and to apply an adaptive similarity threshold. Therefore, relying solely on $\phi(\cdot,\cdot)$ may be insufficient to capture this complexity. Thus, to explicitly capture confidence trends and address Claim 5.2, we define scalar-valued confidence trend features for each node $v_{i}\in\mathcal{V}$:
$\tau^{(0)}_{i}:=~{}\mathrm{Conf}(v_{i}),\quad\quad\tau^{(k)}_{i}:=~{}{\sum}_{v_{j}\in\widehat{\mathcal{N}}^{(1)}(v_{i})}\widetilde{A}_{i,j}\tau^{(k-1)}_{i}\quad(k\geq 1),$ (3)
@@
where $\widetilde{\mathbf{A}}:=\mathbf{D}^{-0.5}\mathbf{A}\mathbf{D}^{-0.5}$ is the normalized adjacency matrix and $\widetilde{A}_{i,j}$ denotes its $(i,j)$-th entry.
@@
We define the confidence difference between orders as $\Delta\tau_{i}^{(k)}:=\tau_{i}^{(k)}-\tau_{i}^{(k-1)}$ for $k\geq 1$. Based on Claim 5.2, the sign of these differences (e.g., between zeroth and first order, and first and second order) serves as a useful signal for identifying nodes that are endpoints of unlearned edges. Thus, we define the following binary-valued trend feature for each node $v_{i}$:
$\widetilde{\boldsymbol{\tau}}_{i}:=$ $~{}\left[\boldsymbol{1}{\{\Delta\tau_{i}^{(1)}<0\}},\boldsymbol{1}{\{\Delta\tau_{i}^{(1)}>0\}},\boldsymbol{1}{\{\Delta\tau_{i}^{(2)}<0\}},\boldsymbol{1}{\{\Delta\tau_{i}^{(2)}>0\}}\right].$ (4)
The final attack model is defined as:
$g(v_{i},v_{j}):=~{}\sigma(\underbrace{\phi([\mathbf{x}_{i}~{}||~{}\mathbf{p}_{i}],[\mathbf{x}_{j}~{}||~{}\mathbf{p}_{j}])}_{\mathrm{MIA}}+\underbrace{\mathbf{h}^{\top}[\widetilde{\boldsymbol{\tau}}_{i}~{}||~{}\widetilde{\boldsymbol{\tau}_{j}}]}_{\mathrm{Trend}}).$ (5)
@@
In the equation above, the MIA term estimates membership from similarity in features and probabilities, and the trend term compensates for the similarity gap by distinguishing node types as indicated by Claim 5.1. The sigmoid function $\sigma(\cdot)$ maps the output to $[0,1]$, with 0 indicating non-member and 1 indicating member. This design allows the attack model to incorporate an adaptive threshold based on node-specific trend features, while leveraging any existing MIA framework as its base.
@@
Unlearning-Aware Attack Model Training. We train the attack model $g$ on the shadow dataset $\mathcal{G}^{\mathrm{sha}}$ using the outputs of the shadow victim model $f^{\prime}$. Given the shadow query set $\mathcal{Q}_{\mathrm{sha}}$ with known membership labels, we optimize a link-prediction loss:
$\mathcal{L}_{\mathrm{attack}}(\mathcal{Q}_{\mathrm{sha}}):=-{\sum}_{(v_{i},v_{j})\in\mathcal{Q}_{\mathrm{sha}}^{+}}\log g(v_{i},v_{j})-{\sum}_{(v_{i},v_{j})\in\mathcal{Q}_{\mathrm{sha}}^{-}}\log\bigl{(}1-g(v_{i},v_{j})\bigr{)}.$ (6)
@@
Performing TrendAttack. After training, we transfer $g$ to the target unlearned graph $\mathcal{G}_{\mathrm{un}}$ by computing, for each query pair $(v_{i},v_{j})\in\mathcal{Q}$, the feature-probability similarity and trend features from the real model output $f_{\mathcal{G}_{\mathrm{un}}}(\cdot;\boldsymbol{\theta}_{\mathrm{un}})$ and partial graph knowledge $\mathcal{K}$. We then evaluate $g(v_{i},v_{j})$ on predicting whether $(v_{i},v_{j})\in\mathcal{E}_{\mathrm{orig}}$. By combining both similarity and trend signals learned on the shadow graph, $g$ effectively supports unlearning inversion on the real unlearned graph.
6 Experiments
@@
In this section, we describe our experimental setup and present the main empirical results.
6.1 Experiment Settings
@@
Datasets. We evaluate our attack method on four standard graph ML benchmark datasets: Cora yang2016revisiting , Citeseer yang2016revisiting , Pubmed yang2016revisiting , and LastFM-Asia yang2016revisiting . To construct the shadow dataset $\mathcal{G}_{\mathrm{sha}}$ and the real attack dataset $\mathcal{G}_{\mathrm{orig}}$, which should share similar distributions, we use METIS to partition the entire training graph into two balanced subgraphs, one for shadow and one for attack. Following the setting in GIF wu2023gif , we use 90% of the nodes in each subgraph for training and the remaining for testing. All edges between the two subgraphs are removed, and there are no shared nodes, simulating real-world scenarios where shadow and attack datasets are disconnected.
@@
Victim Model. We adopt a two-layer GCN kipf2017semi as the victim model, trained for node classification. Our GCN implementation follows the standard settings used in GIF wu2023gif for consistency and reproducibility. After training the GCN on both the shadow and attack datasets, we perform edge unlearning on 5% of randomly selected edges using standard graph unlearning methods, including GIF wu2023gif , CEU wu2023certified , and Gradient Ascent (GA) wu2024graphguard . We follow the official settings from each method’s paper and codebase to ensure faithful reproduction.
@@
Baselines. We do not compare with unlearning inversion attacks hu2024learn that require access to pre-unlearning models, which is unrealistic under our setting. To demonstrate that unlearning inversion cannot be solved by naive link prediction, we evaluate a simple GraphSAGE model and a state-of-the-art link prediction method, NCN wang2024neural . For membership inference attacks (MIAs) under the same black-box assumption as ours, we consider three widely used methods: StealLink he2021stealing , MIA-GNN olatunji2021membership , and GroupAttack zhang2023demystifying . We follow their official hyperparameter settings from their respective papers and repositories. All experiments are run five times, and we report the mean and the standard error.
Evaluation Metrics. We evaluate our attack on the attack dataset using a specific query set $\mathcal{Q}$. We randomly select 5% of the edges as unlearned edges $\mathcal{Q}^{+}_{\mathrm{un}}$ (label 1), and another 5% of remaining edges as regular member edges $\mathcal{Q}^{+}_{\mathrm{mem}}$ (label 1). We then sample an equal number of non-existent (negative) edges $\mathcal{Q}^{-}$ (label 0) such that $|\mathcal{Q}^{-}|=|\mathcal{Q}^{+}_{\mathrm{un}}|+|\mathcal{Q}^{+}_{\mathrm{mem}}|$. We use AUC as the primary evaluation metric. Since both unlearned edges and regular member edges are important for membership inference, we compute the overall AUC on the full query set $\mathcal{Q}=\mathcal{Q}^{+}_{\mathrm{un}}\cup\mathcal{Q}^{+}_{\mathrm{mem}}\cup\mathcal{Q}^{-}$.
@@
More experimental details, including model parameters, baselines, and datasets, are in Appendix F.
6.2 Comparison Experiments
@@
Table 1: Main Comparison Results. We present the AUC scores for attack methods across different edge groups. The best results are highlighted in bold, while the second-best results are underlined.
In this study, we present a comprehensive comparison of all our baselines mentioned in Section 6.1, and the results are shown in Table 1. Specifically, to demonstrate that our method better captures unlearned edges, we evaluate AUC across three groups: Unlearned ($\mathcal{Q}^{+}_{\mathrm{un}}\cup\mathcal{Q}^{-}$), Original ($\mathcal{Q}^{+}_{\mathrm{mem}}\cup\mathcal{Q}^{-}$), and All (the entire $\mathcal{Q}$). A small gap between Unlearned and Original AUCs indicates a better balance in the model’s predictions. The full table with standard deviation can be found in Appendix G.
We consider two variants of our method, TrendAttack-MIA and TrendAttack-SL, which adopt MIA-GNN olatunji2021membership and StealLink he2021stealing as their respective backbone models. From the table, we make the following observations: (i) Compared with their MIA prototypes, both variants of TrendAttack significantly improve the gap between Unlearned and Original AUCs, as well as the overall AUC. This demonstrates the effectiveness of our proposed trend-based attack and unlearning-aware training design. (ii) Overall, our proposed attack, especially TrendAttack-SL, achieves the best performance on most datasets and unlearning methods. The only failure case is on Citeseer + CEU + Original, where the gap to StealLink is small. This does not undermine our contribution, as our primary focus is on the unlearned sets, and this case is a special exception. TrendAttack-MIA achieves the second-best results across many settings, while its relatively lower performance is mainly due to the weak backbone model (MIA-GNN), not the attack framework itself. (iii) Among all baselines, attack-based methods consistently outperform link prediction methods, showing that link prediction alone cannot effectively solve the unlearning inversion problem. Among the attack baselines, StealLink is the strongest, while MIA-GNN performs poorly as it only uses output probabilities and ignores features.
6.3 Ablation Studies
@@
Figure 3: Ablation study on the impact of victim models.
Impact of Victim Models. From Table 1, we observe that the proposed attack remains stable against unlearning methods. In this study, we further investigate whether TrendAttack’s performance maintains stability with respect to changes in victim models. Specifically, we fix the unlearning method to GIF and use TrendAttack-SL as our model variant. We compare it against the best-performing baseline, StealLink, with results shown in Figure 3. From the figure, we see that our proposed model consistently outperforms the baseline, demonstrating stability across different victim models. An interesting observation is that performance significantly improves on GAT compared to other baselines, suggesting that GAT may be more vulnerable to model attacks.
@@
(b) Pubmed results.
(c) LastFM-Asia results.
Figure 4: Ablation study on the impact of unlearning ratio.
Due to space limitations, more experiments can be found in Appendix G.
7 Conclusion
In this work, we study a novel and challenging problem of graph unlearning inversion attack, which aims to recover unlearned edges using black-box GNN outputs and partial knowledge of the unlearned graph. To tackle this problem, we identify two key intuitions, probability gap and confidence pitfall, which motivate a simple yet effective attack framework, TrendAttack, revealing a potential privacy risk of current graph unlearning methods. Despite the effectiveness of TrendAttack, it has several limitations. First, our method focuses solely on link-level inference, and does not consider to node-level or feature-level membership inference. Second, while we provide empirical evidence for the probability gap intuition in the appendix, we do not offer a formal theoretical justification. These limitations highlight important directions for future research.
Appendix

### v2

@@ 5.2.1. Attack Model
It is well established that the edge existence between $v_{i}$ and $v_{j}$ can be inferred from the similarity between their features and output probabilities (olatunji2021membership; he2021stealing). Since prior GNN privacy attacks have developed a variety of similarity computation frameworks for this task, we adopt a general formulation that computes a scalar similarity score between $v_{i}$ and $v_{j}$ as $\phi([\mathbf{x}_{i}~||~\mathbf{p}_{i}],[\mathbf{x}_{j}~||~\mathbf{p}_{j}])$, where $\phi$ is an arbitrary similarity function. This formulation is flexible and covers several existing MIA methods. For example, when $\phi([\mathbf{x}_{i}~||~\mathbf{p}_{i}],[\mathbf{x}_{j}~||~\mathbf{p}_{j}])=\mathbf{h}^{\top}\cdot\mathrm{MLP}(\mathbf{p}_{i},\mathbf{p}_{j})$, the model recovers MIA-GNN (olatunji2021membership). It can also recover the StealLink attack (he2021stealing) by incorporating their manually defined similarity features (e.g., cosine similarity, JS divergence, etc.) into $\phi$. This flexible structure allows our attack model to incorporate any existing MIA method as the backbone, ensuring it performs at least as well as prior approaches.
@@
In addition, as indicated by Claim 5.1 and Claim 5.3, a key challenge in graph unlearning inversion is to distinguish nodes associated with unlearned edges from others and to apply an adaptive similarity threshold. Therefore, relying solely on $\phi(\cdot,\cdot)$ may be insufficient to capture this complexity. Thus, to explicitly capture confidence trends and address Claim 5.3, we define scalar-valued confidence trend features for each node $v_{i}\in\mathcal{V}$ as follows:
$\tau^{(0)}_{i}$:= $~\max_{\ell}\mathbf{p}_{i,\ell},$
(3) $\tau^{(l)}_{i}$:= $~\sum_{v_{j}\in\widehat{\mathcal{N}}^{(1)}(v_{i})}\widetilde{A}_{i,j}\tau^{(l-1)}_{i}\quad(l=1,2,\ldots,k),$
@@
where $k$ is the maximum order, $\widetilde{\mathbf{A}}:=\mathbf{D}^{-0.5}\mathbf{A}\mathbf{D}^{-0.5}$ is the normalized adjacency matrix, and $\widetilde{A}_{i,j}$ is its $(i,j)$-th entry.
Here, the zeroth-order feature $\tau^{(0)}_{i}$ is the model confidence of $f_{\mathcal{G}_{\mathrm{un}}}(\cdot;\bm{\theta}_{\mathrm{un}})$ for node $v_{i}$, based on its output probability $\mathbf{p}_{i}$. Higher-order features $\tau^{(l)}_{i}$ are obtained by aggregating lower-order features from the neighborhood of $v_{i}$. Note that computing $\widetilde{A}_{i,j}$ may not necessarily require the full adjacency matrix $\mathbf{A}$, since the normalization can be done using only the nodes accessible to the attacker.
@@
Next, we define the confidence difference between orders as $\Delta\tau_{i}^{(k)}:=\tau_{i}^{(k)}-\tau_{i}^{(k-1)}$ for $k\geq 1$. Based on Claim 5.3, the sign of these differences (e.g., between zeroth and first order, and first and second order) serves as a useful signal for identifying nodes that are endpoints of unlearned edges. Thus, we define the following binary-valued trend feature for each node $v_{i}$:
(4) $\widetilde{\bm{\tau}}_{i}^{(k)}:=\mathop{\Big\|}_{l=1}^{k}\left[\bm{1}\{\Delta\tau_{i}^{(l)}<0\},\bm{1}\{\Delta\tau_{i}^{(l)}>0\}\right],$
where $\|$ denotes vector concatenation, $\bm{1}\{\cdot\}$ denotes the indicator function, and the resulting $k$-th order trend feature $\widetilde{\bm{\tau}}_{i}^{(k)}$ captures whether the confidence difference is positive or negative at each hop of neighbor for node $v_{i}$.
In practice, we find that using only the first-order trend feature (i.e., $\widetilde{\bm{\tau}}_{i}^{(1)}$) is sufficient to achieve superior attack performance (see Figure 4 in Supplementary G). This means we only need the black-box model output for the first-hop neighbors of the node pair of interest $(v_{i},v_{j})$, demonstrating that our method only requires minimal additional knowledge. The final attack model is given as:
(5) $g(v_{i},v_{j}):=~\sigma(\underbrace{\phi([\mathbf{x}_{i}~||~\mathbf{p}_{i}],[\mathbf{x}_{j}~||~\mathbf{p}_{j}])}_{\mathrm{Similarity}}+\underbrace{\mathbf{h}^{\top}[\widetilde{\bm{\tau}}_{i}~||~\widetilde{\bm{\tau}_{j}}]}_{\mathrm{Trend}}).$
@@
In the equation above, the similarity term estimates membership from similarity in features and probabilities, and the trend term compensates for the similarity gap by distinguishing node types as indicated by Claim 5.1. The sigmoid function $\sigma(\cdot)$ maps the output to $[0,1]$, with 0 indicating non-member and 1 indicating member. This design allows the attack model to incorporate an adaptive threshold based on node-specific trend features, while leveraging any existing MIA framework with $\phi(\cdot,\cdot)$ as its base.
5.2.2. Shadow Victim Model Training
To train an attack model that predicts whether $(v_{i},v_{j})\in\mathcal{G}_{\mathrm{orig}}$, we need ground-truth labels to optimize its parameters. To obtain these labels, we simulate the training and unlearning process of the target GNN by creating a shadow dataset $\mathcal{G}^{\mathrm{sha}}$. We first train a shadow victim model on this dataset and then unlearn a subset of edges, allowing us to generate the necessary ground-truth labels for training the attack model. Below, we provide details of the shadow and attack models.
To construct an attack model $g(v_{i},v_{j})$ that accurately predicts membership information given partial knowledge of the unlearned graph $\mathcal{K}_{\mathcal{Q}}:=(\mathcal{P}_{\mathcal{Q}},\mathcal{F}_{\mathcal{Q}})$, we first simulate the victim model’s behavior using a shadow dataset $\mathcal{G}^{\mathrm{sha}}$ and a shadow victim GNN $f^{\prime}_{\mathcal{G}_{\mathrm{un}}^{\mathrm{sha}}}(\cdot;\bm{\theta}_{\mathrm{un}}^{\prime})$. This model is trained on a node classification task and then unlearns a small subset of edges $\Delta\mathcal{G}^{\mathrm{sha}}$. Specifically, the pre- and post-unlearning parameters are obtained similarly via Eq. (1) and Eq. (3) in Section 3.
This training setup explicitly models unlearning behavior, in contrast to prior MIA approaches that rely solely on the original model $f^{\prime}_{\mathcal{G}^{\mathrm{sha}}_{\mathrm{orig}}}(\cdot;\bm{\theta}_{\mathrm{orig}}^{\prime})$ and do not account for the effects of unlearning. Using this shadow victim model, we construct a shadow query set $\mathcal{Q}_{\mathrm{sha}}$ and its associated partial knowledge $\mathcal{K}_{\mathcal{Q}_{\mathrm{sha}}}:=(\mathcal{P}_{\mathcal{Q}_{\mathrm{sha}}},\mathcal{F}_{\mathcal{Q}_{\mathrm{sha}}})$ from the features, connectivity, and outputs of $f^{\prime}$ on $\mathcal{G}^{\mathrm{sha}}$, which are then used to train the attack model.
5.2.3. Attack Model Training
@@
We train the attack model $g$ on the shadow dataset $\mathcal{G}^{\mathrm{sha}}$ using the outputs of the shadow victim model $f^{\prime}$. Given the shadow query set $\mathcal{Q}_{\mathrm{sha}}$ with known membership labels, we optimize a link‐prediction loss:
(6) $\mathcal{L}_{\mathrm{attack}}(\mathcal{Q}_{\mathrm{sha}}):=\!\!\!\!\!\!\!\sum_{(v_{i},v_{j})\in\mathcal{Q}_{\mathrm{sha}}^{+}}\!\!\!\!\!\!\log g(v_{i},v_{j})-\!\!\!\!\!\!\!\sum_{(v_{i},v_{j})\in\mathcal{Q}_{\mathrm{sha}}^{-}}\!\!\!\!\!\!\log(1-g(v_{i},v_{j})).$
The pseudo-code for building the shadow victim model and training the attack model $g$ on the shadow dataset $\mathcal{G}^{\mathrm{sha}}$ is provided in Algorithm 1 in Supplementary C.
5.2.4. Performing TrendAttack.
@@
After training, we apply the attack model $g$ to the target unlearned graph $\mathcal{G}_{\mathrm{un}}$ by computing, for each query pair $(v_{i},v_{j})\in\mathcal{Q}$, the feature-probability similarity and trend features from the real model output $f_{\mathcal{G}_{\mathrm{un}}}(\cdot;\bm{\theta}_{\mathrm{un}})$ and partial graph knowledge $\mathcal{K}$. We then evaluate $g(v_{i},v_{j})$ on predicting whether $(v_{i},v_{j})\in\mathcal{E}_{\mathrm{orig}}$. By combining both similarity and trend signals learned on the shadow graph, $g$ effectively supports unlearning inversion on the real unlearned graph. The pseudo-code for the attack process is provided in Algorithm 2 in Supplementary C.
6. Experiments
@@
In this section, we describe our experimental setup and present the main empirical results of this work.
6.1. Experiment Settings
@@
Datasets. We evaluate our attack method on four standard graph ML benchmark datasets: Cora (yang2016revisiting), Citeseer (yang2016revisiting), Pubmed (yang2016revisiting), and LastFM-Asia (yang2016revisiting). To construct the shadow dataset $\mathcal{G}_{\mathrm{sha}}$ and the real attack dataset $\mathcal{G}_{\mathrm{orig}}$, which should share similar distributions, we use METIS to partition the entire training graph into two balanced subgraphs, one for shadow and one for attack. Following the setting in GIF (wu2023gif), we use 90% of the nodes in each subgraph for training and the remaining for testing. All edges between the two subgraphs are removed, and there are no shared nodes, simulating real-world scenarios where shadow and attack datasets are disconnected.
@@
Victim Model. We adopt a two-layer GCN (kipf2017semi) as the victim model, trained for node classification. Our GCN implementation follows the standard settings in GIF (wu2023gif) for consistency and reproducibility. After training the GCN on both the shadow and attack datasets, we perform edge unlearning on 5% of randomly selected edges using standard graph unlearning methods, including GIF (wu2023gif), CEU (wu2023certified), and Gradient Ascent (GA) (wu2024graphguard). We follow the official settings from each method’s paper and codebase to ensure faithful reproduction.
@@
Baselines. We do not compare with unlearning inversion attacks (hu2024learn) that require access to pre-unlearning models, which is unrealistic under our setting. To demonstrate that unlearning inversion cannot be solved by naive link prediction, we evaluate a simple GraphSAGE model and a state-of-the-art link prediction method, NCN (wang2024neural). For membership inference attacks (MIAs) under the same black-box assumption as ours, we consider three widely used methods: StealLink (he2021stealing), MIA-GNN (olatunji2021membership), and GroupAttack (zhang2023demystifying). We follow their official hyperparameter settings from their respective papers and repositories. All experiments are run five times, and we report the mean and the standard error.
[…]
@@
More experimental details, including model parameters, baselines, and datasets, are in Supplementary F.
6.2. Unlearning Inversion Attack Performance
Table 1. Main Comparison Results. We present the AUC scores for attack methods across different edge groups. The best results are highlighted in bold, while the second-best results are underlined.
@@ 6.2. Unlearning Inversion Attack Performance
In this study, we present a comprehensive comparison of all our baselines mentioned in Section 6.1, and the results are shown in Table 1. Specifically, to demonstrate that our method better captures unlearned edges, we evaluate AUC across three groups: Unlearned ($\mathcal{Q}^{+}_{\mathrm{un}}\cup\mathcal{Q}^{-}$), Original ($\mathcal{Q}^{+}_{\mathrm{mem}}\cup\mathcal{Q}^{-}$), and All (the entire $\mathcal{Q}$). A small gap between Unlearned and Original AUCs indicates a better balance in the model’s predictions. The full table with standard deviation can be found in Supplementary G.
We consider two variants of our method, TrendAttack-MIA and TrendAttack-SL, which adopt MIA-GNN (olatunji2021membership) and StealLink (he2021stealing) as their respective backbone models. From the table, we observe: (i) Compared with their MIA prototypes, both variants of TrendAttack significantly improve the gap between Unlearned and Original AUCs, as well as the overall AUC. This demonstrates the effectiveness of our proposed trend-based attack and unlearning-aware training design; (ii) Overall, our proposed attack, especially TrendAttack-SL, achieves the best performance on most datasets and unlearning methods. The only failure case is on Citeseer + CEU + Original, where the gap to StealLink is small. This does not undermine our contribution, as our primary focus is on the unlearned sets, and this case is a special exception. TrendAttack-MIA achieves the second-best results across many settings, while its relatively lower performance is mainly due to the weak backbone model (MIA-GNN), not the attack framework itself; and (iii) Among all baselines, attack-based methods consistently outperform link prediction methods, showing that link prediction alone cannot effectively solve the unlearning inversion problem. Among the attack baselines, StealLink is the strongest, while MIA-GNN performs poorly as it only uses output probabilities and ignores features.
6.3. Ablation Studies
@@
Figure 3. Ablation study on the impact of victim models.
Impact of Victim Models. From Table 1, we observe that the proposed attack remains stable against unlearning methods. In this study, we further investigate whether TrendAttack’s performance maintains stability with respect to changes in victim models. Specifically, we fix the unlearning method to GIF and use TrendAttack-SL as our model variant. We compare it against the best-performing baseline, StealLink, with results shown in Figure 3. From the figure, we can find that our proposed model consistently outperforms the baseline, demonstrating stability across different victim models. An interesting observation is that performance significantly improves on GAT compared to other baselines, suggesting that GAT may be more vulnerable to model attacks.
@@
(b) Citeseer results.
(c) Pubmed results.
Figure 4. Ablation study on the impact of trend feature order. Overall attack AUC as a function of the trend feature order (0–3) for three unlearn methods across three datasets.
Impact of Trend Feature Orders. We also study the influence of trend feature orders on the performance of the proposed TrendAttack. Following the experimental setup described in Section 6.2, we employ the most effective variant, TrendAttack-SL, for this ablation study. Specifically, we vary the trend feature order $k$ (see Algorithm 2) to assess the trade-off between incorporating additional neighborhood information and achieving high attack performance. The results are presented in Figure 4, from which we make the following observations: (i) Incorporating trend features of orders 1, 2, or 3 significantly improves attack performance compared to the 0-th order (which corresponds to a degenerate form of TrendAttack that reduces to a simple StealLink attack). This highlights the effectiveness of our proposed trend feature design; and (ii) The attack performance remains relatively stable across orders 1 to 3, indicating that the method is robust to the choice of trend order. Notably, lower-order features (e.g., order 1) still have strong performance while requiring less auxiliary neighborhood information, making them more practical in real-world inversion attack scenarios.
Impact of Edge Unlearning Ratio. Due to space limitations, more experiments on the impact of edge unlearning ratio can be found in Figure 8 in Supplementary G.
(a) Cora results.
(b) Citeseer results.
(c) Pubmed results.
Figure 5. Transferability of attack models across different GNN architectures on Cora, Citeseer, and Pubmed. Each heatmap shows attack AUC when the shadow model (rows) and attack model (columns) differ.
6.4. Transferability of TrendAttack
Model Transferability. We study the transferability of attack models across different GNN architectures. Specifically, we evaluate how attacks trained on one shadow victim model generalize to a different target victim model. We use GIF as the victim unlearning method and select TrendAttack-SL, our strongest model variant, for evaluation. Figures 5 (a)-(c) show the results on Cora, Citeseer, and Pubmed, respectively.
In each figure, the diagonal entries correspond to cases where the shadow and target models are the same, while off-diagonal entries represent transferability experiments where the shadow and target models differ. We make two key observations: (i) Our attack is robust to differences between shadow and target models. Most transfer results show only minor performance drops. The largest drop occurs for GAT on Citeseer, where the AUC decreases from 0.9175 to 0.8601 when the shadow model is GCN, which remains acceptable. (ii) Choosing a strong shadow victim model can slightly improve attack performance. As shown in Figure 3, GAT has the highest attack AUC. Using SGC or GCN to attack GAT results in a performance drop, while using GAT to attack other models can improve results. Therefore, selecting a shadow model with high attack performance can slightly boost transferability.
Dataset Transferability. Due to space limitations, we defer the additional dataset transferability results to Figure 9 in Supplementary G.
7. Conclusion
In this work, we address the novel and challenging problem of graph unlearning inversion attacks, aiming to recover unlearned edges using black-box GNN outputs and partial knowledge of the unlearned graph. We identify two key intuitions, probability similarity gap and confidence pitfall, which inspire a simple yet effective attack framework, TrendAttack, revealing potential privacy risks in current graph unlearning methods. Several promising directions remain. While our study focuses on link-level inversion, extending TrendAttack to node- or feature-level attacks would be valuable. Moreover, although the probability similarity gap intuition is supported empirically, a formal theoretical justification could further strengthen the understanding of graph unlearning inversion. We hope these future directions can further advance this early study of privacy risks in graph unlearning and provide valuable insights.
Acknowledgments
This material is based upon work supported by, or in part by, the Army Research Office (ARO) under grant number W911NF-2110198, the Department of Homeland Security (DHS) under grant number 17STCIN00001-05-00, and Cisco Faculty Research Award. Xiaorui Liu is supported by the National Science Foundation (NSF) Award under grant number IIS-2443182. The findings in this paper do not necessarily reflect the view of the funding agencies.
Supplementary Material